Explain uncertainty principles.


1
Expert's answer
2021-04-23T11:01:49-0400

Answer

According to uncertainty principle position and velocity of and object can not be measured exactly at same time.

In mathematical form it can be given as

"\Delta x\Delta p\geq \frac{h}{4\pi}"